If you copy a page on a machine at 68​%, you should get a similar copy of the page. What is the corresponding setting to obtain the original from the​ copy?

In a proper fraction.
Answer #1

If you want to increase the 68% to a percentage which will give the original.

Therefore 0.68* x = 1

or x = 1/0.68

= 100/68 % (In Proper Fraction)

= 147.05%

If you find the answer helpful, please give a upvote/like/thumbs up. Thank you in advance. It means alot to us!!